# Leveraging Sub-Optimal Data for Human-in-the-Loop Reinforcement Learning

### v1

@@ A.2 Training details
Ensemble size 3
Frequency of feedback
20000 (DMControl)
10000 (Window-open, Door-unlock)
5000 (Hammer, Drawer-open, Door-open, Door-lock)
Feedback budget for ablations
2000 (Door-open)
100 (Cheetah-run, Walker-walk)
# of feedback queries per session
50 (Hammer, Drawer-open, Door-open, Door-lock)
20 (All DMControl tasks, Window-open, Door-unlock)
10 (Cheetah-run scalar fb experiment)
Sampling scheme
Uniform (DMControl)
Disagreement Sampling (Meta-world)
# of training updates 50
# of NN layers 4 (including output layer)
Intermediate NN activations Leaky relu
Final output activation
Leaky ReLU (PEBBLE and Deep TAMER)
Leaky ReLU (SDP all except Tanh for Hammer,
Drawer-open, Walker-walk (preference feedback exp).)
Hidden units 128
Loss
Cross entropy (Preference learning experiments)
MSE (Scalar feedback experiments)
Optimizer Adam
@@
Appendix B Additional analysis
Figure 5 emphasizes how the agent update phase does result in new transitions, therefore the teacher provides feedback to transitions that are different from the original sub-optimal transitions used for pre-training. Figure 6 shows the true reward values for sub-optimal data gathered through a random policy in the DMControl suite. This emphasizes that the true reward value is close to the value we pseudo-label the sub-optimal transitions with (i.e., zero), therefore SDP should not yield a large incorrect reward bias.
@@ Appendix B Additional analysis
Figure 7: These figures show the reward model weights of SDP after the reward model pre-training phase. This demonstrates that the reward model pre-train phase of SDP does not result in zero neural network weights.
Figure 7 demonstrates that the reward model pre-training phase does not produce reward model weights of zero, emphasizing why we do not experience the same performance degradation that occurs if we use a zero-weight initialization.
(a) SDP phase ablation: Cheetah-run, preference feedback
(b) SDP phase ablation: Door-open, preference feedback
(c) Fake input ablation: Cheetah-run, preference feedback
Figure 8: SDP phase and fake input studies: Figures 8a and 8b demonstrate the importance of using both components of SDP, reward pre-training (green) and agent update (grey). Figure 8c compares SDP pre-trained with real sub-optimal transitions to SDP pre-trained with fake transitions.
Figures 8a and 8b show additional phase ablations in the preference learning experiments done on the Cheetah-run and Door-open tasks. This highlights the importance of using both phases in SDP. Furthermore, Figure 8c shows another example of SDP when pre-trained with fake transitions in Cheetah-run. This emphasizes that SDP achieves its best performance when pre-trained with real sub-optimal transitions an RL agent experiences.
(a) Amount of prior data study: Cheetah-run, preference feedback
(b) Amount of prior data study: Door-open, preference feedback
(c) Amount of feedback study: Door-open, preference feedback
Figure 9: Amount of pre-training data and feedback studies: Figures 9a and 9b compare the number of sub-optimal data used in SDP. Figure 9c compares the number of feedback queries and their effect on SDP and PEBBLE.
@@
Figures 9a and 9b show additional ablations over the amount of prior data used in SDP and Figure 9c ablates the number of feedback queries. Overall, we observed similar performance gains as described in section 5.3.
(a) Relabeling sub-opt data study: Walker-walk, scalar feedback
(b) Using high-quality data in SDP study: Cheetah-run, preference feedback
Figure 10: Figure 10a shows the results of not relabeling sub-optimal data transitions with the updated reward model. Figure 10b shows the results of pseudo-labeling high-quality data transitions with zeros in SDP.
@@
SDP is combined with three preference learning algorithms, PEBBLE, RUNE, and SURF. A core feature of these algorithms is that every time the reward model is updated, all transitions inside the RL agent’s replay buffer are updated using the latest reward model. Figure 10a shows the effects of not relabeling the sub-optimal data (i.e., the reward labels) with the latest learned reward model. This means the reward labels remain frozen at zero throughout the entire training process. This ablation was to demonstrate that if the sub-optimal data in the agent’s replay buffer is not updated with the latest reward model, then performance will suffer. This is likely the case because the incorrect reward bias from the pseudo-labeling process persists, whereas when we relabel the transitions with the updated reward model, the incorrect reward bias may reduce over time.
@@
Moreover, we show that the effectiveness of SDP relies on the use of sub-optimal data transitions. If we use high-quality data transitions (i.e., transitions that came from a fully trained RL agent policy), SDP will fail (see Figure 10b). This unsurprising result confirms that pseudo-labeling high-reward transitions with zero can significantly hurt the reward model and the agent’s performance
[…]
Figure 11: Imperfect feedback study: These figures compare the effect of noisy teachers on SDP and PEBBLE (or R-PEBBLE).
@@
Tables 4 and 5 provide a summary of the mean final performance and the mean area under the curve for all environments and benchmarks.
Task - Feedback Metric R-PEBBLE SDP + R-PEBBLE Deep TAMER SAC
Walker-walk - 100
Final Perf.
AUC
362.99 ± 181.22*
13147.1 ± 6102.31*
931.31 ± 36.59
34981.65 ± 1559.24
33.10 ± 11.10*
1823.42 ± 578.48*
961.46 ± 6.21
41973.52 ± 593.89
Cheetah-run - 60
Final Perf.
AUC
522.10 ± 238.87*
33557.58 ± 15185.47*
862.72 ± 49.13
55144.89 ± 4266.94
30.71 ± 16.26*
2459.23 ± 1226.34*
794.24 ± 43.93
53996.18 ± 2814.56
Quadruped-walk - 200
Final Perf.
AUC
543.92 ± 193.21*
37076.52 ± 11968.48*
777.38 ± 156.74
58370.85 ± 10759.0
73.74 ± 48.38*
7365.68 ± 3556.27*
912.98 ± 33.84
58110.69 ± 12486.76
@@
Table 4: This table shows the performance (mean +/- 95% confidence intervals) for all scalar feedback experiments.
* indicates that SDP + the base preference learning algorithm achieves a statistically greater score (p $<$ 0.05) than the base preference learning algorithm alone (i.e., without SDP).
Task - Feedback Metric PEBBLE SDP + PEBBLE RUNE SDP + RUNE SURF SDP + SURF SAC
Walker-walk - 100
Final Perf.
AUC
168.38 ± 118.07*
7224.14 ± 3184.56*
487.24 ± 49.71
17939.94 ± 1412.96
163.05 ± 72.50*
6232.93 ± 2724.95*
575.35 ± 121.51
20614.29 ± 3543.59
211.39 ± 109.20*
7334.85 ± 3132.28*
420.19 ± 49.16
16288.09 ± 1588.21
961.46 ± 6.21
41973.52 ± 593.89
Cheetah-run - 200
Final Perf.
AUC
579.35 ± 74.58*
38966.57 ± 5818.25*
735.31 ± 95.76
51342.85 ± 3772.22
677.57 ± 82.51
45538.51 ± 5134.30
622.58 ± 317.32
36646.1 ± 17981.84
645.63 ± 90.88
40234.13 ± 3388.71
648.27 ± 128.68
42323.53 ± 4395.56
794.24 ± 43.93
53996.18 ± 2814.56
Quadruped-walk - 500
Final Perf.
AUC
373.12 ± 282.63*
20834.8 ± 11054.15*
741.91 ± 105.44
42074.4 ± 12172.93
377.91 ± 261.50*
22780.57 ± 10608.41
700.03 ± 111.22
32454.73 ± 4036.69
599.04 ± 208.38
31076.33 ± 12776.06
764.71 ± 156.54
44115.58 ± 9547.31
912.98 ± 33.84
58110.69 ± 12486.76
Door-unlock - 500
Final Perf.
AUC
40.0 ± 37.70
2110.0 ± 1852.85*
70.0 ± 29.07
4400.0 ± 1170.56
24.0 ± 37.50
1196.0 ± 1476.11
28.0 ± 34.17
1876.0 ± 2378.14
18.0 ± 35.28*
1498.0 ± 2617.88*
72.0 ± 20.93
4650.0 ± 1711.70
100.0 ± 0.0
8594.0 ± 135.74
Window open - 200
Final Perf.
AUC
28.0 ± 35.82
732.0 ± 1093.36
54.0 ± 33.72
1656.0 ± 922.40
10.0 ± 19.60
218.0 ± 341.87
28.0 ± 21.82
1024.0 ± 847.59
8.0 ± 9.60*
202.0 ± 252.33*
70.0 ± 35.61
2018.0 ± 1005.70
100.0 ± 0.0
4432.0 ± 150.22
Drawer open - 1000
Final Perf.
AUC
0.0 ± 0.0
348.0 ± 642.97
26.0 ± 32.56
1622.0 ± 1891.32
0.0 ± 0.0*
42.0 ± 50.51*
64.0 ± 18.18
3860.0 ± 1597.17
20.0 ± 39.20*
714.0 ± 1374.95*
76.0 ± 14.67
4072.0 ± 1916.90
100.0 ± 0.0
8068.0 ± 844.38
Door open - 2000
Final Perf.
AUC
80.0 ± 39.2
5908.0 ± 2894.91
98.0 ± 3.92
6510.0 ± 897.07
80.0 ± 39.20
5570.0 ± 2809.87
56.0 ± 45.38
3424.0 ± 2587.34
100 ± 0.0
7282.0 ± 490.13
86.0 ± 22.85
6366.0 ± 714.27
100.0 ± 0.0
9020.0 ± 254.72
Door-lock - 500
Final Perf.
AUC
34.0 ± 28.81*
788.0 ± 676.40*
90.0 ± 8.77
2274.0 ± 261.83
58.0 ± 29.98
1266.0 ± 662.60
40.0 ± 37.19
918.0 ± 717.44
76.0 ± 37.50
1846.0 ± 910.79
80.0 ± 12.40
1932.0 ± 224.79
96.0 ± 4.8
3718.0 ± 110.25
Hammer - 7500
Final Perf.
AUC
2.0 ± 3.92*
462.0 ± 236.13*
74.0 ± 17.09
2122.0 ± 1104.20
16.0 ± 31.36*
466.0 ± 288.1*
68.0 ± 29.98
2496.0 ± 684.39
0.0 ± 0.0*
536.0 ± 244.93*
52.0 ± 34.73
3066.0 ± 1085.67
100.0 ± 0.0
4538.0 ±2550.30
Table 5: This table shows the performance (mean +/- 95% confidence intervals) for all preference learning experiments.
* indicates that SDP + the base preference learning algorithm achieves a statistically greater score (p $<$ 0.05) than the base preference learning algorithm alone (i.e., without SDP).

### v2

@@ A.2 Training Details
Last Layer Neural Network Activation Leaky ReLU
Table 5: Tuned hyperparameters for human feedback experiments.
Appendix B Human Subject Study Details
Our study was approved by an external ethics committee. In total, 16 users participated in our study, 3 of which participated in both Pendulum-swingup and Cartpole-swingup tasks. As for the amount of participation per task, 7 users provided preferences in Pendulum-swingup and 12 users provided preferences in Cartpole-swingup. During the timeline of our user study, we initially had participants interacting in Pendulum-swingup, however we found that it was very simple, therefore we decided to include Cartpole-swingup, as a more difficult task. Each participant provided preferences for a single run of the environment, as each session took approximately 1 to 1.5 hours. Table 5 shows the specific hyperparameters used.
B.1 Protocol Steps
The study took place in person, where the participant was first provided the consent from to review and sign. Then, participants and the researcher, together, reviewed the instructions outlining the objective of the study. The instructions had two components. First, they outlined participant’s goal in the study. Second, as done in Christiano et al. (2017), we provided a guide on what constitutes good and bad behavior in each domain.
Pendulum-swingup
The first set of instructions for the study are as follows:
  1. 1.
    Objective: for the pole to swing up and balance.
  2. 2.
    Agent controls how much torque (or force or twist) to apply
  3. 3.
    It does not matter which way the pole swings, left or right, as long as the pole balances.
  4. 4.
    Your task: You will see two clips of behaviors, and your job is to select the clip you think is better given the above objective. If you think both clips are identical in behavior, you can select equally preferable.
The second set of instructions (i.e., advice) for the study are as follows:
  1. 1.
    The first priority is for the pole to begin swinging back and forth (i.e., picking up momentum). Therefore, the video clip where the pole has swung higher is better. Even if the agent is not behaving well in either clips, if you can tell that the pole is higher in one clip than the other, it is better to prefer that clip.
  2. 2.
    In general, if the pole is swinging rapidly in a circle (e.g., complete 360), then this is usually worse behavior than if the pole is barely moving.
Cartpole-swingup
The first set of instructions for the study are as follows:
  1. 1.
    Pole is attached to a cart.
  2. 2.
    Agent can move the cart left and right.
  3. 3.
    Goal is for the agent to move the cart such that the pole swings up and balances
  4. 4.
    Your task: You will see two clips of behaviors, and your job is to select the clip you think is better given the above objective. If you think both clips are identical in behavior, you can select equally preferable.
The second set of instructions (i.e., advice) for the study are as follows:
  1. 1.
    The first priority is for the pole to begin swinging back and forth (i.e., picking up momentum). Therefore, the video clip where the pole has swung higher is better. Even if the agent is not behaving well in either clips, if you can tell that the pole is higher in one clip than the other, it is better to prefer that clip.
  2. 2.
    In general, if the pole is swinging rapidly in a circle (e.g., complete 360), then this is usually worse behavior than if the pole is barely moving.
  3. 3.
    If the pole is balancing then falls over, then is is better behavior than if the pole is barely moving.
During the instruction period, participants also watched video clips of what successful and unsuccessful behavior looks like. In addition, participants were able to ask questions about the environment’s objective.
Once the instruction period was complete, participants completed a practice round in providing preferences. This was included so users could gain familiarity with the interface. Figure 7 shows a screenshot of the interface used. Participants would use keyboard input to move from one video clip to the next. Then, participants had the opportunity to rewatch either clips as many times as they liked. Afterwards, participants chose which clip they preferred (or equally preferred) via keyboard input. Each video clip consisted of a 50 step segment, which was approximately 2 seconds long. The total study time was $\sim(1-1.5)$ hours. After the study was complete, participants filled out a demographic survey which included questions pertaining to their age, race/ethnicity, education level completed, and area of expertise.
Figure 7: This shows the user interface used for the human subject study. Participants would view each video clip, sequentially, then decide which clip they preferred.
B.2 Preference Learning Survey Details
Table 6 provides a list of all papers included in our preference learning survey. We now describe the criteria for inclusion and exclusion in our survey. To find papers, we used Google Search with key words: preference learning, reinforcement learning from human feedback, and preference based reinforcement learning. We also found papers by reviewing those that cited popular preference learning works. This included Christiano et al. (2017) and Lee et al. (2021a), as these were two of the first papers that popularized preference learning. We included conference, journal and workshop papers in our survey. We did not include papers that were strictly on Arxiv. To keep the survey within scope, we did not include any preference learning works related to large language or foundation models. We also did not include any works that were entirely theoretical (e.g., no empirical results were provided).
Table 6: Articles used in Preference Learning Survey from Section 5.3. NA indicates no information was provided.
Paper Reference Human Users (Non-Author) Number of Users
(Myers et al., 2023) Yes 22
(Metcalf et al., 2024) Yes 50
(Myers et al., 2022) Yes 50
(Christiano et al., 2017) Yes NA
(Hejna III & Sadigh, 2023) Yes 4
(Hwang et al., 2023b) Yes 5
(Biyik & Sadigh, 2018) Yes 10
(Hwang et al., 2023a) Yes 5
(Bıyık et al., 2020) Yes 10
(Wilde et al., 2021) Yes 18
(Bıyık et al., 2022) Yes 15
(Ibarz et al., 2018) Yes NA
(Knox et al., 2023) Yes 143
(Holk et al., 2024) Yes 32
(Metcalf et al., 2023) Yes 40
(Marta et al., 2024) Yes NA
(Marta et al., 2023a) Yes 70
(Marta et al., 2023b) Yes 20
(Ren et al., 2022) Yes NA
(Mehta & Losey, 2024) Yes 15
(White et al., 2024) Yes 20
(Wang et al., 2022) Yes 10
(Wang et al., 2023) No
(Zhang & Kashima, 2024) No
(Lee et al., 2021a) No
(Liang et al., 2022) No
(Park et al., 2022) No
(Hu et al., 2024) No
(Liu et al., 2022) No
(Liu et al., 2023) No
(Xue et al., 2024) No
(Daniels-Koch & Freedman, 2022) No
(Verma et al., 2023) No
(Barnett et al., 2023) No
(Verma & Metcalf, 2024) No
(Metcalf et al., 2023) No
(Swamy et al., 2024) No
(Lee et al., 2021b) No
(Wang et al., 2021) No
(Maxence Hussonnois & Rana, 2023) No
(Wilson et al., 2012) No
(Liu et al., 2024) No
(Giovanelli et al., 2024) No
(Zhu et al., 2024) No
(Cheng et al., 2024) No
Appendix C Additional Human Teacher Results
Task Feedback Background AUC P Value Final Performance P Value
Cartpole-swingup 48
Non-CS
CS
8258.44 ± 1898.47
10208.11 ± 2311.16
0.183
648.88 ± 35.17
614.14 ± 114.81
0.692
Pendulum-swingup 40
Non-CS
CS
7677.84 ± 3540.70
10188.70 ± 3858.94
0.258
327.65 ± 407.80
641.68 ± 237.76
0.23
Table 7: This table shows the AUC and final performance of SDP (mean ± 95% confidence intervals) for CS and non-CS participants in the human subject study.
Task Feedback Demographic AUC P Value
Cartpole-swingup 48 Female 8720.03 ± 3185.4 0.085
Male 11770.22 ± 1669.96
Pendulum-Swingup 40 Female 8587.13 ± 2772.96 0.331
Male 10134.67 ± 4822.22
Cartpole-swingup 48 Age 18-30 10109.54 ± 2370.61 0.313
Age 50-70 10923.05 ± 1292.83
Pendulum-Swingup 40 Age 18-30 8619.07 ± 3381.48 0.318
Age 50-70 10607.93 ± 5249.72
Cartpole-swingup 48 Education: Bachelors or higher 11044.55 ± 1783.25 0.794
Education: High school diploma 6247.99 ± 5186.47
Pendulum-Swingup 40 Education: Bachelors or higher 10188.88 ± 3858.95 0.742
Education: High school diploma 7677.85 ± 3540.7
Table 8: This table shows the AUC of SDP (mean ± 95% confidence intervals) for different demographic conditions. This table highlights that SDP can be effective irrespective of demographic background (gender, age, and educational background).
Tables 7-8 highlight that there is no significant difference in the performance of SDP when human teachers have differing backgrounds. This includes a background in CS, educational level, age, and gender. We note that we did not perform statistical analysis comparing racial groups as some experiments had only one participant of a particular racial background, making comparisons less meaningful.
Appendix D Additional Simulated Teacher Results
D.1 Study of Agent Update Phase
Figure 8 emphasizes how the agent update phase does result in new transitions, therefore the teacher provides feedback to transitions that are different from the original sub-optimal transitions used for pre-training.
@@ D.1 Study of Agent Update Phase
D.2 True Reward Values of Random Policy
Figure 9 shows the true reward values for sub-optimal data gathered through a random policy in the DMControl suite. This emphasizes that the true reward value is close to the value we pseudo-label the sub-optimal transitions with (i.e., zero), therefore SDP should not yield a large incorrect reward bias.
@@ D.2 True Reward Values of Random Policy
D.3 Zero Weight Study
To understand the effect of each component of SDP, we performed two additional ablations. First, in the reward model pre-training phase, the goal is for the reward model to learn to output zero. However, a trivial means to achieve an output of zero is to set all weights and biases in the neural network to zero. Therefore, we compare the full SDP to SDP using a zero-weight initialization as a replacement for the reward model pre-training. We found that using a zero-weight initialization for the reward model in place of the pre-training phase results in significantly degraded performance (see the green curve in Figure 10–right). This is not surprising, as previous works have found that a zero-weight initialization can negatively affect the training of neural networks (Blumenfeld et al., 2020; Zhao et al., 2022). In addition, Figure 11 demonstrates that the reward model pre-training phase does not produce reward model weights of zero, emphasizing why we do not experience the same performance degradation that occurs if we use a zero-weight initialization.
D.4 Fake Input Study
[…]
Figure 10: Zero weights and fake input studies in Walker-walk
@@ D.4 Fake Input Study
Figure 11: This figure shows the reward model weights of SDP after the reward model pre-training phase. This demonstrates that the reward model pre-train phase of SDP does not result in zero neural network weights.
Figure 12: Fake Input Study in Cheetah Run Preference Learning
D.5 SDP Component Studies
Figure 13 shows additional phase ablations in the preference learning experiments done on the Cheetah-run environment. This highlights the importance of using both phases in SDP. In addition, we ran an experiment in Walker-walk where we directly apply the procedure Yu et al. (2022) uses for leveraging sub-optimal data in the offline RL setting. They simply store the sub-optimal transitions in the RL agent’s replay buffer with the pseudo reward label of 0 and follow standard offline RL. We repeated this with the other difference of following standard preference learning (PEBBLE) afterwards. We found that the average final performance was 104.32 with a 95% confidence interval of 46.12, which is close to 4 times less than the final performance of SDP.
Figure 13: Phase Ablation Study in Cheetah Run Preference Learning
@@ D.5 SDP Component Studies
Figure 14 shows an additional ablation over the amount of prior data used in SDP. We observed similar performance gains as described in section 5.4.
Figure 14: Prior Data Amount Study in Cheetah Run Preference Learning
D.6 Effect of Relabelling Replay Buffer
@@
SDP is combined with four preference learning algorithms, PEBBLE, RUNE, SURF, and MRN. A core feature of these algorithms is that every time the reward model is updated, all transitions inside the RL agent’s replay buffer are updated using the latest reward model. Figure 15 shows the effects of not relabeling the sub-optimal data (i.e., the reward labels) with the latest learned reward model. This means the reward labels remain frozen at zero throughout the entire training process. This ablation was to demonstrate that if the sub-optimal data in the agent’s replay buffer is not updated with the latest reward model, then performance will suffer. This is likely the case because the incorrect reward bias from the pseudo-labeling process persists, whereas when we relabel the transitions with the updated reward model, the incorrect reward bias may reduce over time.
Figure 15: Relabeling sub-opt data study: Walker-walk, scalar feedback
D.7 Effect of Transition Quality in SDP
@@
Moreover, we show that the effectiveness of SDP relies on the use of sub-optimal data transitions. If we use high-quality data transitions (i.e., transitions that came from a fully trained RL agent policy), SDP will fail (see Figure 16). This unsurprising result confirms that pseudo-labeling high-reward transitions with zero can significantly hurt the reward model and the agent’s performance
Figure 16: Using high-quality data in SDP study: Cheetah-run, preference feedback
D.8 Scalar-based Experiment Statistics
@@
Tables 9 and 10 provide a summary of the mean final performance and the mean area under the curve for all environments and benchmarks in the scalar feedback setting
Learning from Scalar Feedback
Task SDP R-PEBBLE Deep TAMER
Walker-walk 931.31 ± 36.59 362.99 ± 181.23* 33.10 ± 11.10*
Cheetah-run 862.72 ± 49.13 522.10 ± 238.87* 30.71 ± 16.26*
Quadruped-walk 777.38 ± 156.74 543.92 ± 193.21* 73.74 ± 48.38*
Table 9: This table shows the mean final performance plus and minus the 95% confidence interval. * indicates SDP achieves a significantly greater mean final performance.
Learning from Scalar Feedback
Task SDP R-PEBBLE Deep TAMER
Walker-walk 34981.65 ± 1559.24 13147.1 ± 6102.31* 1823.42 ± 578.48*
Cheetah-run 55144.89 ± 4266.94 33557.58 ± 15185.47* 2459.23 ± 1226.34*
Quadruped-walk 58370.85 ± 10759.00 37076.52 ± 11968.48* 7365.68 ± 3556.27*
Table 10: This table shows the mean area under the learning curve (AUC) plus and minus the 95% confidence interval. * indicates SDP achieves a significantly greater AUC.
D.9 Preference Learning Experiment Statistics
Tables 11-14 provide a summary of the mean final performance and the mean area under the curve for all environments and benchmarks in the preference feedback setting.
Task Feedback Method Final Return P Value
Cartpole-swingup 48 PEBBLE 226.4 ± 136.35 0.007*
SDP + PEBBLE 533.94 ± 57.71
RUNE 263.48 ± 176.2
SDP + RUNE 561.77 ± 108.17 0.021*
SURF 83.12 ± 44.53
SDP + SURF 324.19 ± 179.8 0.039*
MRN 151.05 ± 76.31
SDP + MRN 644.03 ± 57.99 0.0*
Walker-walk 100 PEBBLE 173.9 ± 100.26 0.001*
SDP + PEBBLE 490.07 ± 50.07
RUNE 160.6 ± 61.85 0.0*
SDP + RUNE 575.71 ± 104.52
SURF 202.1 ± 95.17 0.005*
SDP + SURF 435.99 ± 39.45
MRN 215.14 ± 58.09 0.0*
SDP + MRN 672.3 ± 85.53
Cheetah-run 200 PEBBLE 582.51 ± 67.62 0.062
SDP + PEBBLE 704.63 ± 102.27
RUNE 681.59 ± 68.05 0.645
SDP + RUNE 616.78 ± 279.34
SURF 642.52 ± 84.3 0.266
SDP + SURF 682.34 ± 65.08
MRN 728.36 ± 38.69 0.305
SDP + MRN 748.99 ± 55.59
Quadruped-walk 500 PEBBLE 350.24 ± 228.91 0.017*
SDP + PEBBLE 753.18 ± 104.79
RUNE 363.65 ± 163.93 0.009*
SDP + RUNE 704.1 ± 102.16
SURF 550.51 ± 211.19 0.089
SDP + SURF 767.78 ± 140.67
MRN 200.12 ± 63.04 0.0*
SDP + MRN 733.83 ± 62.87
@@ D.9 Preference Learning Experiment Statistics
Table 11: This table shows the final performance (mean ± 95% confidence intervals) for all DMControl preference learning experiments. * indicates significant differences.
Task Feedback Method AUC P Value
Cartpole-swingup 48 PEBBLE 8248.14 ± 2148.59 0.0*
SDP + PEBBLE 21519.9 ± 2435.02
RUNE 8817.21 ± 3534.57 0.004*
SDP + RUNE 20970.54 ± 4643.96
SURF 5593.44 ± 1099.56 0.074
SDP + SURF 10288.22 ± 4584.16
MRN 6774.95 ± 1783.95 0.0*
SDP + MRN 22564.32 ± 3846.6
Walker-walk 100 PEBBLE 7571.79 ± 2903.09 0.001*
SDP + PEBBLE 18458.81 ± 1224.69
RUNE 6525.52 ± 2566.88 0.0*
SDP + RUNE 21059.76 ± 3056.28
SURF 7675.24 ± 2857.99 0.001*
SDP + SURF 16700.51 ± 1449.62
MRN 7711.05 ± 1295.46 0.0*
SDP + MRN 25184.26 ± 3071.65
Cheetah-run 200 PEBBLE 39245.94 ± 5189.16 0.006*
SDP + PEBBLE 51121.68 ± 3488.87
RUNE 45824.63 ± 4459.04 0.792
SDP + RUNE 37202.31 ± 16330.83
SURF 40522.18 ± 2989.22 0.218
SDP + SURF 42777.1 ± 3775.07
MRN 47569.45 ± 2191.57 0.106
SDP + MRN 49968.24 ± 2187.97
Quadruped-walk 500 PEBBLE 20584.95 ± 9814.49 0.018*
SDP + PEBBLE 42044.34 ± 11183.84
RUNE 22861.89 ± 9861.31 0.082
SDP + RUNE 32664.79 ± 3929.32
SURF 30884.8 ± 11660.52 0.074
SDP + SURF $4$ 8404.76
MRN 14805.07 ± 1860.39 0.002*
SDP + MRN 35209.32 ± 6225.09
Table 12: This table shows the AUC (mean ± 95% confidence intervals) for all DMControl preference learning experiments. * indicates significant differences.
Task Feedback Method Final Return P Value
Hammer 7500 PEBBLE 10.0 ± 13.58 0.003*
SDP + PEBBLE 66.0 ± 21.18
RUNE 4.0 ± 7.01 0.001*
SDP + RUNE 68.0 ± 17.0
SURF 0.0 ± 0.0 0.033*
SDP + SURF 36.0 ± 25.16
MRN 4.0 ± 7.01 0.027*
SDP + MRN 46.0 ± 27.5
Door-unlock 500 PEBBLE 42.0 ± 34.35 0.066
SDP + PEBBLE 80.0 ± 15.68
RUNE 26.0 ± 22.59 0.02*
SDP + RUNE 36.0 ± 38.65
SURF 20.0 ± 35.06 0.354
SDP + SURF 80.0 ± 22.17
MRN 48.0 ± 38.17 0.399
SDP + MRN 56.0 ± 37.02
Door-lock 500 PEBBLE 38.0 ± 30.06 0.035*
SDP + PEBBLE 80.0 ± 7.84
RUNE 62.0 ± 32.51 0.793
SDP + RUNE 40.0 ± 30.87
SURF 70.0 ± 31.85 0.684
SDP + SURF 60.0 ± 13.58
MRN 66.0 ± 26.93 0.268
SDP + MRN 78.0 ± 17.88
Drawer-open 1000 PEBBLE 4.0 ± 7.01 0.045*
SDP + PEBBLE 36.0 ± 25.16
RUNE 0.0 ± 0.0 0.001*
SDP + RUNE 66.0 ± 14.24
SURF 20.0 ± 35.06 0.018*
SDP + SURF 80.0 ± 14.67
MRN 20.0 ± 35.06 0.136
SDP + MRN 56.0 ± 40.21
Window-open 200 PEBBLE 34.0 ± 37.44 0.234
SDP + PEBBLE 54.0 ± 26.35
RUNE 12.0 ± 21.04 0.014*
SDP + RUNE 38.0 ± 24.42
SURF 14.0 ± 20.44 0.098
SDP + SURF 66.0 ± 26.35
MRN 46.0 ± 32.61 0.208
SDP + MRN 68.0 ± 31.06
Table 13: This table shows the final performance (mean ± 95% confidence intervals) for all Metaworld preference learning experiments. * indicates significant differences.
Task Feedback Method AUC P Value
Hammer 7500 PEBBLE 424.0 ± 175.18 0.017*
SDP + PEBBLE $2$.0 ± 1009.47
RUNE 490.0 ± 300.85 0.001*
SDP + RUNE 2520.0 ± 627.47
SURF 482.0 ± 215.59 0.006*
SDP + SURF 3042.0 ± 1073.54
MRN 494.0 ± 311.76 0.016*
SDP + MRN 2494.0 ± 1108.52
Door-unlock 500 PEBBLE 2142.0 ± 1738.06 0.031*
SDP + PEBBLE 4838.0 ± 1251.13
RUNE 1272.0 ± 1275.71 0.047*
SDP + RUNE 1836.0 ± 2063.68
SURF 1580.0 ± 2462.83 0.348
SDP + SURF 4802.0 ± 1537.81
MRN 3082.0 ± 2232.43 0.342
SDP + MRN 3842.0 ± 2230.38
Door-lock 500 PEBBLE 876.0 ± 647.87 0.008*
SDP + PEBBLE 2300.0 ± 135.0
RUNE 1342.0 ± 590.05 0.809
SDP + RUNE 884.0 ± 633.87
SURF 1892.0 ± 842.85 0.429
SDP + SURF 1986.0 ± 184.83
MRN 1966.0 ± 751.51 0.306
SDP + MRN 2236.0 ± 479.9
Drawer-open 1000 PEBBLE 314.0 ± 533.02 0.117
SDP + PEBBLE 1702.0 ± 1706.12
RUNE 38.0 ± 49.71 0.004*
SDP + RUNE 3794.0 ± 1306.6
SURF 696.0 ± 1167.88 0.015*
SDP + SURF 4078.0 ± 1839.07
MRN 1550.0 ± 2690.98 0.211
SDP + MRN 3206.0 ± 2114.16
Window-open 200 PEBBLE 794.0 ± 1015.04 0.131
SDP + PEBBLE 1698.0 ± 822.8
RUNE 230.0 ± 334.14 0.01*
SDP + RUNE 948.0 ± 631.92
SURF 242.0 ± 248.44 0.064
SDP + SURF 2048.0 ± 867.93
MRN 1252.0 ± 796.8 0.223
SDP + MRN 1728.0 ± 666.48
Table 14: This table shows the area under the curve (mean ± 95% confidence intervals) for all Metaworld preference learning experiments. * indicates significant differences.